OpenStudy (anonymous):

Please i need help with my math math is not a good subject for me (determine whether each number is a solution of the given inequality 8m-6≤10 A.2 B.0 C.5

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Maybe A?

OpenStudy (anonymous):

Yea sorry i dont know much about this math is my worst subject:(

OpenStudy (skullpatrol):

No problem, it takes a lot of practice...so replace "m" in your inequality with "2" and simplify the left hand side...do you get a true statement in the end?

OpenStudy (skullpatrol):

Did I ask you to do that?

OpenStudy (skullpatrol):

replace "m" in your inequality with "2" and simplify the left hand side.

OpenStudy (anonymous):

wouldnt it be B then?

OpenStudy (skullpatrol):

There are actually two answers to this question.

OpenStudy (anonymous):

but wouldnt it just be B.?

OpenStudy (skullpatrol):

Yes, but look at A also...

OpenStudy (skullpatrol):

The concept of a variable is very important to understand in math, it means you can REPLACE the letter with a numeral, right?

OpenStudy (anonymous):

Right

OpenStudy (skullpatrol):

Is 10 less than or equal to 10?

OpenStudy (anonymous):

equal?

OpenStudy (skullpatrol):

10 = 10

OpenStudy (anonymous):

K

OpenStudy (skullpatrol):

So, 10 less than or equal to 10 is a true statement and that makes it a solution.

OpenStudy (anonymous):

I just told u it was equal..

OpenStudy (skullpatrol):

I am looking at the truth of the statement to find out if it is a solution.

OpenStudy (anonymous):

Ok..

OpenStudy (skullpatrol):

B. says -6≤10 which is obviously true.

OpenStudy (anonymous):

Right

OpenStudy (skullpatrol):

variable: A symbol used to represent one or more numbers.

OpenStudy (skullpatrol):

Replacing each variable in an open sentence by each of the values in its domain is a way to find solutions of an open sentence.

OpenStudy (anonymous):

I have another one that seems harder then first one do u think u can help me one more time..?

OpenStudy (skullpatrol):

I can try...

OpenStudy (anonymous):

thank u i really appreciate your help

OpenStudy (anonymous):

\[6-n \n{ \ge }{ 11 }\]

OpenStudy (anonymous):

that n should be under the 6-n

OpenStudy (anonymous):

options are A.0.5 B.2 C.4

OpenStudy (skullpatrol):

$$\frac{6-n}{n} \ge 11$$

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

What does option A. mean?

OpenStudy (anonymous):

Honestly I don't know

OpenStudy (skullpatrol):

What do letters stand for?

OpenStudy (anonymous):

A letter in math can stand for any number

OpenStudy (skullpatrol):

Yes, so replace the letter "n" with option A.

OpenStudy (anonymous):

both n's ?

OpenStudy (skullpatrol):

$$\frac{6-n}{n} \ge 11$$ A. 0.5 $$\frac{6-0.5}{0.5} \ge 11$$

OpenStudy (anonymous):

okay i did that

OpenStudy (skullpatrol):

Yes, both n's...now simplify the left hand side

OpenStudy (anonymous):

i came out with \[11\ge11\]

OpenStudy (skullpatrol):

Is that statement true or false?

OpenStudy (anonymous):

False.

OpenStudy (skullpatrol):

Why is it false that 11 (is greater than) OR (equal to) 11

OpenStudy (anonymous):

Well its not greater then because there the same number

OpenStudy (anonymous):

Someone please help me

OpenStudy (skullpatrol):

They are equal, so that makes the statement true.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

wait soo it would be A.?

OpenStudy (skullpatrol):

That is one solution. You still have to check the others...

OpenStudy (anonymous):

Oh okaay

OpenStudy (anonymous):

the next one came out as \[4\ge11\]

OpenStudy (skullpatrol):

True or false

OpenStudy (anonymous):

false.?

OpenStudy (skullpatrol):

$$\frac{6-n}{n} \ge 11$$ let n=2 $$\frac{6-2}{2} \ge 11$$

OpenStudy (skullpatrol):

$$\frac{4}{2} \ge 11$$

OpenStudy (skullpatrol):

Is 2 greater than or equal to 11?

OpenStudy (anonymous):

greater..?

OpenStudy (skullpatrol):

2 is greater than 11?

OpenStudy (anonymous):

i mean equal sorry

OpenStudy (skullpatrol):

2 = 11?

OpenStudy (anonymous):

Idk im lost like very lost thats why i need help

OpenStudy (skullpatrol):

$$\frac{6-n}{n} \ge 11$$ We start with this inequality and we want to find "What value" will make it true?

OpenStudy (anonymous):

the value that completes the problem..

OpenStudy (skullpatrol):

Yes, if by "completes" you mean "gives you a TRUE statement"?

OpenStudy (anonymous):

Yes

OpenStudy (skullpatrol):

If we let n=2 we get $$\frac{6-n}{n} \ge 11$$ $$\frac{6-2}{2} \ge 11$$ $$\frac{4}{2} \ge 11$$$$2 \ge 11$$ Is this true?

OpenStudy (skullpatrol):

Yes or no?

OpenStudy (anonymous):

Yes

OpenStudy (skullpatrol):

Why?

OpenStudy (skullpatrol):

Why is $$2\ge11$$true?

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):